# Gravitational Wave Signal Denoising and Merger Time Prediction By Deep Neural Network

### v1

@@ Abstract
The mergers of massive black hole binaries could generate rich electromagnetic emissions, which allow us to probe the environments surrounding these massive black holes and gain deeper insights into the high energy astrophysics. However, due to the short timescale of binary mergers, it is crucial to predict the time of the merger in advance to devise detailed observational plans. The significant noise and the slow accumulation of signal-to-noise ratio in the inspiral phase make this task particularly challenging. To address this issue, we propose a novel deep neural denoising network in this study, capable of denoising continuous inspiral phase signals lasting up to 30 days. Following the denoising process, we perform the detection and merger time prediction based on the denoised signals. Our results demonstrate that for inspiral phase data with a signal-to-noise ratio between 10 and 50 occurring no more than 10 days before the merger, our prediction error for the merger time is generally within 24 hours.
@@ I INTRODUCTION
In 2015, Adv-LIGO achieved the first direct detection of gravitational waves (GW) [1], marking the beginning of a new era in astronomy. Led by LIGO-Virgo collaboration, nearly a hundred GW events have been discovered [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], greatly enriching our understanding of the universe. However, due to terrestrial noise, ground-based GW detectors typically exhibit sensitivity to frequencies above 10 Hz [15, 16, 17], restricting the scope of astronomical exploration. Consequently, initiatives have been taken to establish space-based observation systems, such as the Laser Interferometer Space Antenna (LISA) [18, 19], Taiji [20, 21], and Tianqin [22], to circumvent the limitations and broaden the observable frequency range. Space-based GW detectors are expected to be able to detect GW signals in the range of 0.1mHz - 1Hz. Within this frequency range, a diverse array of astrophysical sources is anticipated to be detected, including massive black hole binaries (MBHBs), stellar-mass black hole binaries, extreme mass-ratio inspirals, galactic double white dwarfs, and the stochastic gravitational wave background. These detections are expected to provide novel insights and enhance the understanding of cosmic phenomena.
@@
Among these signals, coalescing MBHBs, with signal-to-noise ratios (SNRs) ranging from tens to several thousand [18, 23, 24], are considered the most promising GW sources detectable and thus are designated as the primary science goals for space-based GW detection missions. The merger of MBHBs is anticipated to along with a high black hole accretion rate and a high star formation rate in galaxy [25, 26], which could potentially result in strong electromagnetic (EM) emissions. The EM emission from MBHBs is expected to manifest in various forms during the inspiral, merger, and ring-down phases. During the inspiral phase, the predominant form of EM emission is X-rays [27], which is anticipated to transition to ultraviolet radiation as the merger approaches [28]. In the merger phase, it is expected that a wide variety of EM emissions will be produced, including super-Eddington flares [29], X-rays [30], and jets caused by the surrounding magnetic field [31]. During the ringdown phase, the EM emission is expected to persist for an extended period and predominantly manifest itself as infrared radiation [32]. Observations of EM emission provide the possibility to gain insights into the environment surrounding massive black holes. In recent years, substantial research has concentrated on GW and EM multi-messenger observations of MBHBs [33, 24, 34, 35]. These observations are poised to enhance our understanding of the co-evolution of massive black holes, nuclear star clusters, and their host galaxies [36]. Moreover, the concurrent analysis of GW and EM signals will further our comprehension of the evolution of galaxies at high redshifts [37]. However, the observable duration of some EM emissions is short [38, 39, 40], presenting new challenges for multi-messenger observations of mergers of MBHBs. Consequently, it is necessary to conduct pre-merger detection, i.e., predicting the time of the merger, in order to devise detailed and timely observation plans. This enhances the feasibility of achieving multi-messenger observations for mergers of MBHBs.
@@
In principle, the matched filtering algorithm, which is the predominant method for detecting GW signals, yields the most precise results [41, 42, 43]. However, to obtain the most matched template, an extensive template bank is required [44, 45, 46]. Cornish et al. [47] introduced a low-latency detection pipeline based on matched filtering that can analyze several months of LISA data within a few hours. Additionally, for Tianqin data, Chen et al. [48] designed a search algorithm based on matched filtering by reducing the number of frequency points computed and selecting simpler likelihood functions for probability assessment. However, this also requires approximately 10 hours to execute the search. Given that a large number of sources are likely to be detected in the future, it is essential to develop faster and more accurate search algorithms to further reduce computational complexity and resource consumption.
@@
Deep learning algorithms have demonstrated remarkable efficacy across various GW data analysis tasks due to their robust feature representation capabilities and rapid inference proficiency. These applications include signal detection [49, 50, 51, 52, 53], parameter estimation [54, 55], and signal extraction [56, 57, 58, 59, 60, 61]. Specifically, for the pre-merger detection of MBHB signals, Ruan et al. recently proposed a deep learning model, which enables the detection of GW signals several hours to days before the merger [62]. This development provides the possibility of the simultaneous GW and EM detection of MBHBs. However, this model sets a stringent detection threshold of 0.999 for identifying GW signals. While this threshold significantly reduces false positives, it markedly decreases the model’s sensitivity to GW signals. Additionally, the model does not provide the estimations of the time when the mergers take place, which will be the key parameters for the planning for the detailed planning of observations for EM emissions. The difficulty of pre-merger detection largely stems from the way how the SNR of GW signals from coalescing MBHB are accumulated with time [24]. Especially, the signal from the final hour before the merger contains 99% of the total SNR, causing signals from the inspiral phase to be submerged in overwhelming noise. Consequently, achieving a detectable SNR during the inspiral phase requires substantial time accumulation [18], and the increased length of the signal significantly burdens the training of deep learning models [63]. Most deep learning models designed for GW data analysis can handle data points on the order of . In the mHz frequency band, this corresponds to only a few days’ worth of data with sampling rate . This limitation substantially restricts the capability of deep learning to perform pre-merger detection effectively. Based on the issues discussed above, in this work, we introduce a new model, WaveUformer-CNN. WaveUformer is utilized to denoise 30-day data from the inspiral phase, while the CNN is tasked with predicting the time of the merger events. The design of WaveUformer integrates the effective dimensionality reduction capabilities of WaveUnet [64] with the noise separation properties of Sepformer [65], enabling it to rapidly and efficiently denoise GW signals. Considering the enhanced prominence of features in the denoised signals, the predictive network is intentionally simplistic to prevent overfitting, comprising only a simple CNN architecture.
@@
Our model is capable of denoising 30 days of inspiral phase signals and predicting the corresponding merger time within 0.01 seconds. The training data for the model was set with SNRs ranging from 10 to 50, enabling the model to successfully detect signals even at lower SNRs. Our tests encompass signal denoising, signal detection, and merger time prediction. For the denoising test, our model demonstrated excellent denoising performance, effectively recovering the amplitude and phase of the signals. Specifically, 91.98% of the denoised data achieved an overlap 0.99 with the template data, and 90.52% of the denoised data had an SNR difference of 1 compared to the template. For signal detection, we designed the workflow depicted in Fig 4 and explored the optimal thresholds for detecting signals. The results indicate that our workflow efficiently avoids false positives from pure noise while maintaining high sensitivity to GW signals. For test sets with varying chirp masses, our workflow achieved a minimum detection accuracy of 98.64% for GW signals. Regarding merger time prediction, the model’s prediction error decreases as the merger event approaches, and data with larger chirp masses tend to yield smaller errors. Overall, for data with an SNR between 10 and 50 and occurring no more than 10 days before the merger, our prediction error is generally no more than 24 hours.
II Model Architecture
Our model consists of a denoising module and a merger time prediction module. The denoising module integrates the advantages of WaveUnet and Sepformer, enabling it to reduce noise in long-duration GW sequences effectively. As illustrated in Fig 1, our input data, sampled at 0.1 Hz with a length of 259,200, corresponds to a continuous 30-day period. Initially, the data is subjected to whitening and is normalized between -1 and 1 using min-max normalization, the original maximum and minimum values are recorded for subsequent inverse normalization.
Upon entering the network, the data passes through 12 downsampling blocks with a large kernel size of 15, which enables the model to intricately extract latent features from the GW sequence. After each layer of feature extraction, we preserve the feature vectors for gradual reconstruction of the GW sequence in subsequent upsampling blocks. Notably, before each downsampling block, the data undergoes a coarse-grained feature downsampling operation. Specifically, we perform this feature downsampling using a stride of 2 for each feature vector. This strategy ensures that the model preserves sufficient feature accuracy while significantly reducing the feature length—a critical factor in our model’s ability to handle long GW signals. Ultimately, we obtain the encoded representation of the data, , where represents the channel dimension of the final data, and denotes the feature length of each channel.
The data then passes through a separator module. This module does not directly output GW features but instead estimates a mask for the GW features. This approach enhances the model’s sensitivity to the target features [65, 66]. In this separator, first passes through a layer normalization (LN) [67] followed by a linear layer. Subsequently, overlapping segments of size are created along the time axis with a 50% overlap factor, resulting in , where represents the length of each segment. The Sepformer block, the core component of this separator, comprises a composite structure based on the Transformer [68] architecture. In the Sepformer block, the IntraTransformer and InterTransformer are employed to capture intra-block and inter-block relational patterns, respectively. Both transformers share the same structural design. By swapping the positions of and , we obtain representations of intra-block and inter-block data. The structure of this Transformer block closely resembles the one described in [68]. In our implementation, we denote the input to the Transformer as . Initially, a sinusoidal positional encoding is added to the input , resulting in:
For specific implementation details of the positional encoding, please refer to [68]. Subsequently, multiple Transformer layers are applied. Within each Transformer layer , layer normalization (LN) is first applied, followed by multi-head attention (MHA), which can be represented as follows:
Each attention head computes scaled dot-product attention among all elements in the sequence. The data then passes through a feed-forward network (FFN), which can be expressed as follows:
Thus, the entire Transformer block is defined as:
Our model incorporates two Sepformer blocks. Within each Sepformer block, there are eight IntraTransformer blocks and eight InterTransformer blocks. After passing through these Sepformer blocks, the model employs PReLU activation, followed by a linear layer. It then applies the overlap-add scheme described in [69] to combine the chunks and restore to the original dimensions. The final mask is obtained after processing through a FFN and a ReLU activation function. This mask is then multiplied with the dimensionally reduced representation of the features to yield the GW features.
In each upsampling block, we crop and concatenate the features saved from each corresponding layer of the downsampling blocks with the features of the upsampling layer, facilitating the gradual restoration of the sequence length. Throughout this process, the number of channels progressively decreases, resulting in the denoised waveform being obtained at the final layer.
III Data preparations
@@
In this work, the noisy signal serves as the input for the denoising network, while the GW signal is used as the corresponding label. For the merger time prediction network, the output of the denoising network is utilized as the input, and the merger time , is used as the corresponding label.
@@
We will utilize this PSD combination to obtain simulated noise that complies with the Taiji mission.
@@
All data will undergo whitening and then transformed to the time domain using the inverse real Fast Fourier Transform. We select data spanning 30 days within 0 to 10 days preceding the merger. The template is injected with a specific optimal SNR:
@@
We set the SNR of the training set between 10 and 50, and in accordance with the aforementioned settings, we generate a total of 50,000 training samples.
@@ IV Training strategy
We choose the Adam optimizer [79] combined with a warmup [80] and cosine annealing [81] learning rate schedule to optimize our models effectively. The maximum learning rate is set to , and the minimum to . The total training spans 100 epochs, with the first three epochs constituting the warmup period, during which the learning rate increases from to . This warmup ensures stability in the deeper layers of the training. Afterward, the learning rate enters a cosine annealing phase, gradually decreasing back to . This scheduling helps the model escape local optima and aids in optimal convergence. All models are implemented using the Pytorch [82] framework and were trained 10 hours on three NVIDIA A100 40GB GPUs.

### v2

@@ Abstract
The mergers of massive black hole binaries could generate rich electromagnetic emissions, which allow us to probe the environments surrounding these massive black holes and gain deeper insights into the high energy astrophysics. However, due to the short timescale of binary mergers, it is crucial to predict the time of the merger in advance to devise detailed observational plans. The overwhelming noise and the slow accumulation of signal-to-noise ratio in the inspiral phase make this task particularly challenging. To address this issue, we propose a novel deep neural denoising network in this study, capable of denoising a 30-day inspiral phase signal. Following the denoising process, we perform the detection and merger time prediction based on the denoised signals. Our results demonstrate that for a 30-day inspiral phase data with a signal-to-noise ratio between 10 and 50 occurring no more than 10 days before the merger, our absolute prediction error for the merger time is generally within 24 hours.
@@ I INTRODUCTION
In 2015, Advanced LIGO and Advanced Virgo achieved the first direct detection of gravitational waves (GW) [1], marking the beginning of a new era in astronomy. Led by LIGO-Virgo collaboration, nearly a hundred GW events have been discovered [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], greatly enriching our understanding of the universe. However, due to terrestrial noise, ground-based GW detectors typically exhibit sensitivity to frequencies above 10 Hz [15, 18, 19], restricting the scope of astronomical exploration. Consequently, initiatives have been taken to establish space-based observation systems, such as the Laser Interferometer Space Antenna (LISA) [20, 21], Taiji [22, 23], and Tianqin [24], to circumvent the limitations and broaden the observable frequency range. Space-based GW detectors are expected to be able to detect GW signals in the range of 0.1mHz - 1Hz. Within this frequency range, a diverse array of astrophysical sources is anticipated to be detected, including massive black hole binaries (MBHBs), stellar-mass black hole binaries, extreme mass-ratio inspirals, galactic double white dwarfs, and the stochastic gravitational wave background. These detections are expected to provide novel insights and enhance the understanding of cosmic phenomena.
@@
Among these signals, coalescing MBHBs, with signal-to-noise ratios (SNRs) ranging from tens to several thousand [20, 25, 26], are considered the most promising GW sources detectable and thus are designated as the primary science goal for space-based GW detection missions. The merger of MBHBs is anticipated to along with a high black hole accretion rate and a high star formation rate in galaxy [27, 28], which could potentially result in strong electromagnetic (EM) emissions. The EM emission from MBHBs is expected to manifest in various forms during the inspiral [29, 30], merger [31, 32, 33], and ring-down phases [34]. Observations of EM emission provide the possibility to gain insights into the environment surrounding massive black holes. In recent years, substantial research has concentrated on GW and EM multi-messenger observations of MBHBs [35, 26, 36, 37]. These observations are poised to enhance our understanding of the co-evolution of massive black holes, nuclear star clusters, and their host galaxies [38]. Moreover, the concurrent analysis of GW and EM emission will further our comprehension of the evolution of galaxies at high redshifts [39]. The feasibility of this multimessenger observation largely depends on the ability of space-based GW detectors to accurately localize the source, as well as the capability of telescopes to point to the source’s location. A typical binary black hole system can be localized with a median precision of deg2 ( deg2) at 1 month (1 hour) from the merger [26]. Candidate EM telescopes, including the Square Kilometre Array [40] in radio wave, the Large Synoptic Survey Telescope [41] in optical band, and Athena [42] in X-ray, may require localization precisions ranging from 0.1 to 10 deg2. It is expected that for some sources, multimessenger observations of GW and EM emission may be enabled within a few hours to a few days.
@@
However, the observable duration of some EM emissions may only last for several hours or even shorter periods[43, 44, 45, 46], presenting new challenges for multi-messenger observations of mergers of MBHBs. Consequently, it is necessary to conduct pre-merger detection, i.e., predicting the time of the merger, in order to devise detailed and timely observation plans. This enhances the feasibility of achieving multi-messenger observations for mergers of MBHBs. In principle, the matched filtering algorithm, which is the predominant method for detecting GW signals, yields the most precise results [47, 48, 49]. However, to obtain the template that most closely matches the observed GW signal, an extensive template bank is required [50, 51, 52]. In addition, the computational inefficiency of the matched filtering algorithm poses a significant challenge. Cornish et al. [53] introduced a low-latency detection pipeline based on matched filtering that can analyze several months of LISA data within a few hours. Additionally, for Tianqin data, Chen et al. [54] designed a search algorithm based on matched filtering by reducing the number of frequency points computed and selecting simpler likelihood functions for probability assessment. However, this also requires approximately 10 hours to execute the search. Given that a large number of sources are likely to be detected in the future [35], it is essential to develop faster and more accurate search algorithms to further reduce computational complexity and resource consumption.
@@
Deep learning algorithms have demonstrated remarkable efficacy across various GW data analysis tasks due to their robust feature representation capabilities and rapid inference proficiency. These applications include signal detection [55, 56, 57, 58, 59], parameter estimation [60, 61], and signal extraction [62, 63, 64, 65, 66, 67]. Specifically, for the pre-merger detection of MBHB signals, Ruan et al. recently proposed a deep learning model, which enables the detection of GW signals several hours to days before the merger [68]. This development provides the possibility of the simultaneous GW and EM detection of MBHBs. However, this model sets a stringent detection threshold of 0.999 for identifying GW signals. While this threshold significantly reduces false positives, it markedly decreases the model’s sensitivity to GW signals. This is because they trained the model using noisy data directly. However, the SNR of the inspiral phase signal is too low, and the GW features in the data without denoising are too weak, making it difficult for the model to distinguish between pure noise and GW signals. Therefore, further denoising is necessary to improve the distinguishability between noise and GW signals. Additionally, the model does not provide the estimations of the time when the mergers take place, which will be the key parameters for the planning for the detailed planning of observations for EM emissions. The difficulty of pre-merger detection largely stems from the way how the SNR of GW signals from coalescing MBHB are accumulated with time [26]. Especially, the signal from the final hour before the merger contains 99% of the total SNR, causing signals from the inspiral phase to be submerged in overwhelming noise. Consequently, achieving a detectable SNR during the inspiral phase requires substantial time accumulation [20], and the increased length of the signal significantly burdens the training of deep learning models [69]. Based on the issues discussed above, in this work, we design a deep learning-driven workflow for denoising, detecting, and predicting the corresponding merger time of inspiral signals. The core component of this workflow is WaveUformer-CNN. WaveUformer is utilized to denoise 30-day data from the inspiral phase, while the CNN is tasked with predicting the time of the merger events. The design of WaveUformer integrates the effective dimensionality reduction capabilities of WaveUnet [70] with the noise separation properties of Sepformer [71], enabling it to rapidly and efficiently denoise GW signals. Considering the enhanced prominence of features in the denoised signals, the prediction network is intentionally simplistic to prevent overfitting, comprising only a simple CNN architecture.
@@
Our model is capable of denoising 30 days of inspiral phase signals with a SNR between 10 and 50, and predicting the corresponding merger time within 0.01 seconds. Our tests encompass signal denoising, signal detection, and merger time prediction. For the denoising test, our model demonstrated excellent denoising performance, effectively recovering the amplitude and phase of the signals. Specifically, 91.98% of the denoised data achieved an overlap 0.99 with the template data, and 95.82% of the denoised data has an SNR loss 5% compared to the template SNR. For signal detection, we explored the optimal thresholds for detecting signals using our workflow. The results indicate that our workflow efficiently avoids false positives from pure noise while maintaining high sensitivity to GW signals. Our workflow achieved a average detection accuracy of 99.27% for GW signals. Regarding merger time prediction, the model’s prediction error decreases as the merger event approaches, and data with larger chirp masses tend to yield smaller errors. The absolute error in the merger time prediction for 89.1% of the data is 24 hours.
The organization of this paper is as follows. In Sec II, we introduce the data generation process. In Sec III, we present the workflow used in this work, along with the core deep learning model architecture. Sec IV describes the training strategy of the deep learning model. The results obtained from testing the proposed method on the test set are presented in Sec V. Finally, Sec VI provides a summary.
II Data preparations
@@
In this work, the noisy signal serves as the input for the denoising network, while the GW signal is used as the corresponding label. For the merger time prediction network, the output of the denoising network is utilized as the input, and the merger time , is used as the corresponding label. It is important to note that the above data undergoes whitening preprocessing before being used.
@@
We will utilize this PSD combination to obtain simulated noise that complies with the Taiji mission. At the same time, this PSD will also be used for whitening the GW signal.
@@
We select data spanning 30 days within 0 to 10 days preceding the merger. The sampling frequency of the data is 0.1 Hz, so each data will have 259,200 data points. The template is injected with a specific optimal SNR:
@@
The chosen range for the luminosity distance is from to Mpc, based on the ldc2a [81]. We further analyzed the SNR of the 30-day inspiral signal under this setup, and the results show that the SNR can range from to . The large distribution of SNR is unfavorable for the optimization of the deep learning model. Furthermore, compared to the SNR settings in the work of Ruan et al. [68], we aim to further improve the model’s performance at lower SNR values. Based on the above considerations, we set the SNR of the training set between 10 and 50, with each data point representing a 30-day whitened inspiral signal. The data has a sampling frequency of 0.1 Hz, so the length of each data sample is 259,200. In accordance with the aforementioned settings, we generate a total of 50,000 training samples.
III Methodology
In this work, we designed a workflow as shown on the left side of Fig 1 for denoising, detection, and merger time prediction of inspiral signals. This is a deep learning-driven workflow, and the detailed architecture of the deep learning model is presented on the right side of Fig 1. It is worth noting that the event rate of MBHB overlap detections by the Taiji detector is currently under investigation. Therefore, this work assumes a scenario where MBHBs are relatively far apart, which we believe serves as a good starting point for further research.
Our workflow employs a 30-day time window with a 1-hour moving step to process the GW data stream. Data within the current time window is first subjected to whitening processes, then fed into a denoising model to obtain the denoised waveform. Given the moving step, there exists an overlap of 29 days and 23 hours between the current and the previous time windows. We compute the overlap within this overlapping window between the denoised waveforms of the current and previous windows. If the overlap exceeds a predefined threshold, it indicates the detection of a GW signal. The efficacy of this approach is based on the capability of a well-trained denoising deep learning network to effectively extract deep patterns present in GW sequences. This results in the output of similar denoised waveforms and high overlap in the overlapping window of consecutive time windows. In contrast, such patterns are absent in pure noise signals, leading the model’s outputs to tend towards randomness, thereby resulting in low overlap in the overlapping windows. Upon detection of a GW signal, the denoised GW data is then fed into a prediction network to estimate the merger time.
Our denoising model is based on the WaveUNet [70] architecture, with Sepformer [71] employed in the deeper feature space to enhance the model’s expressive power. The input data , where represents the length of the data, is first processed through 12 downsampling blocks, each consisting of 1D convolution, activation functions, and downsampling operations, to progressively extract deep features, resulting in a deep feature matrix , where represents the channel dimension of the final data, and denotes the feature length of each channel. This feature matrix is then passed through the separator to obtain a GW feature mask , which captures the latent GW features within the deep feature matrix. We multiply the deep feature matrix by the GW feature mask to obtain the separated GW feature matrix . During the upsampling process, 12 upsampling blocks, each consisting of 1D convolution, activation functions, and upsampling operations, are employed to progressively reconstruct the data. Notably, during each upsampling block, the upsampled feature matrix is cropped and concatenated with the corresponding downsampled feature matrix, which allows the model to effectively utilize both shallow and deep features. This structure greatly reduces the information loss incurred during downsampling. Finally, after the last upsampling block, we obtain the denoised data .
In our denoising network, the core component is the separator module, which is also a transformer-based structure. The transformer is a widely proven effective architecture for capturing various complex patterns present in the data [68, 67, 71, 82, 83, 84]. The time complexity of the transformer is generally considered to scale quadratically with the length of the data [82]. In this work, the transformer block is applied to the downsampled deep feature matrix, effectively enhancing the model’s expressive power without significantly increasing the model’s time complexity. In this separator, first passes through a layer normalization (LN) [85] followed by a linear layer. Subsequently, as shown in Fig 2, the feature map is divided into overlapping sub-blocks of size . For each chunk, we employ an IntraTransformer to learn the short-term dependencies present within the chunk. Subsequently, the output block data undergoes permutation, where the size of each new block becomes . The data within each new block represents an integration of the same-position data from all previous blocks, thus each new block contains global information. These new blocks are then fed into an InterTransformer to learn the long-term dependencies within the chunks. Finally, we apply overlap-add [86] to restore the feature representation to the size of . The recovered feature representation is then passed through a feed-forward network (FFN) followed by a ReLU activation function, ultimately producing the . It is worth noting that both the IntraTransformer and InterTransformer employ the same structure as the standard Transformer. The names are used solely to distinguish their respective roles in extracting short-term dependencies and long-term dependencies within the sequence.
We denote the input to the Transformer as . Initially, a sinusoidal positional encoding is added to the input , resulting in:
Since the Transformer operates in parallel, it inherently loses the sequential information present in the data. Positional encoding is used to provide the model with a sense of order by encoding the positional information of the sequence and adding it to the sequence. For specific implementation details of the positional encoding, please refer to [82]. Subsequently, multiple Transformer layers are applied. Within each Transformer layer , layer normalization (LN) is first applied, followed by multi-head attention (MHA), which can be represented as follows:
Each attention head computes scaled dot-product attention among all elements in the sequence. The data then passes through a FFN, which can be expressed as follows:
Thus, the entire Transformer block is defined as:
Our denoising model incorporates two Sepformer blocks. Within each Sepformer block, there are eight IntraTransformer blocks and eight InterTransformer blocks.
@@ IV Training strategy
We choose the Adam optimizer [87] combined with a warmup [88] and cosine annealing [89] learning rate schedule to optimize our models effectively. The maximum learning rate is set to , and the minimum to . The total training spans 100 epochs, with the first three epochs constituting the warmup period, during which the learning rate increases from to . This warmup ensures stability in the deeper layers of the training. Afterward, the learning rate enters a cosine annealing phase, gradually decreasing back to . This scheduling provides a smooth optimization space search strategy. In the early stages of cosine annealing, the learning rate is typically set to a large value, which helps in rapidly exploring the optimization space. Later, the learning rate is systematically reduced according to the cosine schedule, leading to a smoother optimization trajectory. Compared to traditional optimization techniques, cosine annealing enables a more comprehensive exploration of the optimization space, thereby helping to escape local optima and aiding in optimal convergence. All models are implemented using the Pytorch [90] framework and were trained 10 hours on three NVIDIA A100 40GB GPUs.